32x + 23y = 10,100
12x + 14y = 5400
substituttion the varible. I selected equation 2 first to solve for y, but couldn't get y by itself. please help...
subtracted 12x from both sides of 12x + 14y = 5400 I got:
14y = 5400-12x....what is my next step?


Found 2 solutions by edjones, stanbon:
Answer by edjones(8007)

divide the equation by 14.

Answer by stanbon(75887)


32x + 23y = 10,100
12x + 14y = 5400
substitute the varible. I selected equation 2 first to solve for y, but couldn't get y by itself. please help...
subtracted 12x from both sides of 12x + 14y = 5400 I got:
14y = 5400-12x....what is my next step?
-------------
y = (2700-6x)/7
--------
Substitute to get:
32x + 23[2700-6x]/7 = 10100
32x + (62100-138x)/7 = 10100
---
Multiply thru by 7 to get:
224x + 62100 - 138x = 7*10100
86x = 8600
x = 100
=============
Substitute to solve for "y":
y = (2700-6*100)/7
y = (2100)/7
y = 300
==============
Cheers,
Stan H.